# Near-Optimal Adaptive Transforms for LLM Quantization

### v1

@@ Abstract
Quantization to low bitwidth is a standard approach for deploying large language models, however, a few extreme weights and activations stretch the dynamic range and reduce the effective resolution of the quantizer. A common mitigation approach is to apply some fixed orthogonal transforms, such as Hadamard matrices, before quantization, which typically reduces the dynamic range. Yet, these transforms ignore the statistics of the data, and their optimality is currently not understood. In this work, we derive, for the first time, closed-form optimal linear blockwise transforms for joint weight-activation quantization using standard data-free quantizers for common numerical formats. Specifically, we provide derivations of the optimal adaptive (data-aware) transforms for round-to-nearest (RTN), AbsMax-scaled block quantizers for both integer and floating-point formats. The resulting construction, which we call WUSH, combines a Hadamard backbone with a data-dependent component based on second-order moments, yielding a non-orthogonal transform that is provably optimal under mild assumptions and remains structured for efficient implementation. Preliminary experimental results show that our approach consistently improves upon the Hadamard transform for common formats.
@@ 1.1 Overview
Quantization of model weights (NEURIPS2022_c3ba4962; frantar2023optq; MLSYS2024_42a452cb) or activations (pmlr-v202-xiao23c; NEURIPS2024_b5b93943) is now a standard tool for shrinking and accelerating large language models (LLMs) (kurtic2025give), making low-precision inference feasible on a wide range of hardware. A central difficulty, however, is that a few extreme “outlier” weights and activations expand the dynamic range and thereby degrade the effective resolution of low-bit representations.
@@
It is known that one way to mitigate these outliers is to apply linear transforms before quantization (jegou2008hamming; jegou2010product); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (cheequip2023; pmlr-v235-tseng24a; NEURIPS2024_b5b93943; liu2025spinquant). For LLMs, Hadamard rotations have been remarkably effective, and blockwise variants aligned with quantization groups have also been shown to be useful in practice. The recent work of egiazarian2025bridginggappromiseperformance shows that the block-wise Hadamard transform offers the best empirical performance for quantization among a large set of existing orthogonal transforms. Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not depend on the statistics of the underlying weights or activations. This raises a natural question: if the Hadamard transform is not data-aware, in what sense can it be considered optimal for quantization?
@@
In this work, we address this question by deriving closed-form optimal linear blockwise transforms for joint weight-activation quantization, which are generally non-orthogonal and adaptive, i.e., data-aware. As opposed to prior methods such as SpinQuant (liu2025spinquant) or FlatQuant (sun2024flatquant), our approach is closed-form and calibration-based, as it does not require training or fine-tuning.
@@
We call our construction WUSH, and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers. The name WUSH is a mnemonic for its composition. For a layer with blockwise weight $\bm{W}_{\left(i\right)}$ and calibration activation $\bm{X}_{\left(i\right)}$ in the $i$-th block, the WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}$ is constructed as
$\bm{T}_{\mathrm{hsuw}}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
where $\bm{W}_{\left(i\right)}^{\prime}$ is the Cholesky decomposition of $\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}=\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}$, the matrices $\bm{U}_{\left(i\right)},\bm{S}_{\left(i\right)},\bm{V}_{\left(i\right)}$ are from the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$, and $\bm{H}$ is a Hadamard matrix. We will define it again in a more exhaustive manner in Eq. (6). WUSH is explicitly adapted to minimize loss in second-order statistics while retaining a structure amenable to efficient implementation and providing a principled, data-aware alternative to heuristic Hadamard-style and blockwise rotations.
1.2 Related Work
@@
Post-training quantization and outliers. Generic post-training quantization schemes, such as RTN (round-to-nearest), OBQ (frantar2022obc), and GPTQ (frantar2023optq), are highly sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values determine the quantization scale. As a result, most quantization levels are spent covering rare extremes instead of the bulk of the distribution, which leads to a significant loss of accuracy.
@@
Non-uniform bitwidths and explicit outlier storage. One common strategy to mitigate outliers is to use non-uniform bitwidths. Methods such as LLM.int8() (NEURIPS2022_c3ba4962), SpQR (dettmers2024spqr), and QUIK (ashkboos-etal-2024-quik) explicitly separate and store outliers in higher precision, while HPTQ (chen2025geometryllmquantizationgptq) uses Huffman encoding to compress the outliers and reduce their storage costs. These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels and additional engineering effort for efficient deployment.
@@
Transform-based outlier mitigation. Another line of work applies transforms to the weights and activations before quantization to reduce the impact of outliers. SmoothQuant (pmlr-v202-xiao23c) and AWQ (MLSYS2024_42a452cb) rescale channels to stabilize the dynamic ranges of weights and activations. QuIP (cheequip2023) introduces an incoherence processing step. QuIP# (pmlr-v235-tseng24a), QTIP (NEURIPS2024_6de2e84b), and QuaRot (NEURIPS2024_b5b93943) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (liu2025spinquant) learns rotations optimized on calibration data. However, their transforms are either heuristic, costly to learn, or primarily designed for weight-only quantization, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
@@
Blockwise transforms and emerging FP formats. The recent FP formats, such as MXFP (ocp_mx_2023) and NVFP (nvidia_blackwell_2024), have motivated blockwise transform schemes. shao2025blockrotationneedmxfp4 indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but egiazarian2025bridginggappromiseperformance highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions. In contrast, our work derives a closed-form, data-aware optimal blockwise transform (WUSH) and analyzes how such a transform interacts with both FP and INT block quantizers.
1.3 Methodology
@@
Problem setup. We formulate the transformed weight-activation quantization problem as follows. Define $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ as the weight of a linear layer with $d_{\mathrm{in}}$ input channels and $d_{\mathrm{out}}$ output channels. Define $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$ as the calibration input activation with $d_{\mathrm{in}}$ embedding channels and $d_{\mathrm{batch}}$ tokens. Define $q\left(\cdot\right)$ as a quantizer that maps a matrix of continuous values to quantized values (we will specify $q$ later). Let $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}$ be the transforms applied to $\bm{W}$ and $\bm{X}$, respectively. For weight-activation quantization, the output activation $\bm{W}^{\top}\bm{X}$ becomes $q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)$. Our goal is to choose $\bm{T}_{\mathrm{W}}$, $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
@@
Block-independent constraint. While the weights can be pre-quantized, the activations must be transformed and quantized dynamically at inference time. To reduce the computational overhead and simplify the problem, we constrain the transforms $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ to be block-diagonal, and the quantizer $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales. We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$ ($d$ is a factor of $d_{\mathrm{in}}$, and $d$ is a power of $2$). Denote
$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right)\quad\text{and}\quad\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right)$ (3)
with $\bm{T}_{\mathrm{W}_{\left(i\right)}},\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathrm{R}^{d\times d}$. We partition $\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ and $\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{X}_{\left(i\right)}\in\mathrm{R}^{d\times d_{\mathrm{batch}}}$. Then, we can express the output without quantization as $\bm{W}^{\top}\bm{X}=\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}$ and the output with weight-activation quantization as
$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$ (4)
Finally, define the blockwise output loss as
@@
We approximate the layerwise loss as $\ell\approx\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise loss $\ell_{\left(i\right)}$ independently.
@@
Optimal block-diagonal transforms. We will prove that there exist closed-form optimal transforms for each block under reasonable assumptions and smooth approximations, which are optimal for FP and asymptotically optimal for INT quantizations. Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy $\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$, respectively. Without loss of generality (any differences will be canceled in later steps), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$. In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moment before the Cholesky decomposition. Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$. Let $\bm{H}\in\mathbb{R}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix where each element is $\pm d^{-\frac{1}{2}}$. The optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ are constructed as
${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top}\quad\text{and}\quad{\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$ (6)
@@
respectively. Note that ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}={\bm{T}_{\mathrm{hsuw}}^{-\top}}_{\left(i\right)}$  [^1] The -$\top$ superscript notation means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$. , which can be easily verified by calculating ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}{\bm{T}_{\mathrm{hsuw}}^{\top}}_{\left(i\right)}=\mathbf{I}$ with $\mathbf{I}$ as the identity matrix.
@@
Remark. The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Eq. (6), which explains why it has been empirically shown to be one of the most effective data-agnostic orthogonal transforms.

### v2

# WUSH: Near-Optimal Adaptive Transforms for LLM Quantization

@@ Abstract
[…]
@@ 1 Introduction
Quantization of model weights (Dettmers et al., 2022; Frantar et al., 2023; Lin et al., 2024) or activations (Xiao et al., 2023; Ashkboos et al., 2024b) is now a standard tool for shrinking and accelerating large language models (LLMs) (Kurtic et al., 2025), making low-precision inference feasible on a wide range of hardware. A central difficulty, however, is that a few extreme “outlier” weights and activations expand the dynamic range and thereby degrade the effective resolution of low-bit representations.
@@
One way to mitigate such outliers is to apply linear transforms before quantization (Jegou et al., 2008; Jégou et al., 2011); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (Chee et al., 2023; Tseng et al., 2024a; Ashkboos et al., 2024b; Liu et al., 2025). For LLMs, Hadamard rotations have been remarkably effective, and blockwise variants aligned with quantization groups have also been shown to be useful in practice. Egiazarian et al. (2025) recently showed that the blockwise Hadamard transform offers the best empirical performance for quantization among a large set of existing transforms. Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not adapt to the statistics of the underlying weights or activations. This raises a natural question: if the Hadamard transform is not data-aware, in what sense can it be considered optimal for quantization?
@@
In this work, we address this question by deriving closed-form optimal linear blockwise transforms for joint weight-activation quantization, which are generally non-orthogonal and adaptive, i.e., data-aware. As opposed to prior methods such as SpinQuant (Liu et al., 2025) or FlatQuant (Sun et al., 2025), our approach is closed-form and calibration-based, as it does not require training or fine-tuning the entire model.
@@
We call our construction WUSH and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers. (The name WUSH is a mnemonic for its composition.) Consider a block of layer weights $\bm{W}_{\left(i\right)}$, with calibration activation $\bm{X}_{\left(i\right)}$. Then, the WUSH transform ${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}$ is:
${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
where $\bm{W}_{\left(i\right)}^{\prime}$ is the Cholesky decomposition of $\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}=\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}$, the matrices $\bm{U}_{\left(i\right)},\bm{S}_{\left(i\right)},\bm{V}_{\left(i\right)}$ are from the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$, and $\bm{H}$ is a Hadamard matrix. We define the transform more formally in Equation˜8.
[…]
2 Related Work
@@
Post-training quantization (PTQ) and outliers. PTQ schemes, such as RTN (round-to-nearest), OBQ (Frantar and Alistarh, 2022), and GPTQ (Frantar et al., 2023), are sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values can determine the quantization scale, leading to high error.
@@
Non-uniform bitwidths and explicit outlier storage. One common strategy to mitigate outliers is to use non-uniform bitwidths. Methods such as LLM.int8() (Dettmers et al., 2022), SpQR (Dettmers et al., 2024), and QUIK (Ashkboos et al., 2024a) explicitly separate and store outliers in higher precision, while HPTQ (Chen et al., 2025) uses Huffman encoding to compress the outliers and reduce their storage costs. These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels.
@@
Transform-based outlier mitigation. Another line of work applies transforms to the weights and activations before quantization to reduce the impact of outliers. SmoothQuant (Xiao et al., 2023) and AWQ (Lin et al., 2024) rescale channels to stabilize the dynamic ranges of weights and activations. QuIP (Chee et al., 2023) introduces an incoherence processing step. QuIP# (Tseng et al., 2024a), QTIP (Tseng et al., 2024b), and QuaRot (Ashkboos et al., 2024b) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (Liu et al., 2025) and FlatQuant (Sun et al., 2025) learn transforms optimized on calibration data. However, their transforms are heuristic and costly to learn, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
@@
Blockwise transforms and emerging FP formats. The recent FP formats, such as MXFP (MX Alliance, 2023) and NVFP (NVIDIA, 2024), have motivated blockwise transform schemes. Shao et al. (2025) indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but Egiazarian et al. (2025) highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions. In contrast, this work (WUSH) derives a closed-form, data-aware optimal blockwise transform and analyzes how such a transform interacts with both FP and INT block quantizers.
3 Methodology
@@
Problem setup. We formulate the transformed weight-activation quantization problem as follows. Define $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ as the weight of a linear layer with $d_{\mathrm{in}}$ input channels and $d_{\mathrm{out}}$ output channels. Define $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$ as the calibration input activation with $d_{\mathrm{in}}$ embedding channels and $d_{\mathrm{batch}}$ tokens. Define $q\left(\cdot\right)$ as a quantizer that maps a matrix of continuous values to quantized values (we will specify $q$ later). Let $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}$ be the transforms applied to $\bm{W}$ and $\bm{X}$, respectively. For weight-activation quantization, the output activation $\bm{W}^{\top}\bm{X}$ becomes $q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)$. Our goal is to choose $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
@@
Block-independent constraint. While the weights can be pre-quantized, the activations must be transformed and quantized dynamically at inference time. To reduce the computational overhead of these operations, it is common (Egiazarian et al., 2025) to constrain $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}$ to be block-diagonal transforms, and $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales. We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$, with $d$ being a factor of $d_{\mathrm{in}}$ and a power of $2$. Denote the block partitions of the matrices as
$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right),$ $\bm{T}_{\mathrm{W}_{\left(i\right)}}\in\mathbb{R}^{d\times d},$ (3)
$\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right),$ $\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathbb{R}^{d\times d},$
$\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right],$ $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}},$
$\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right],$ $\bm{X}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{batch}}}.$
Then, we can express the output without and with weight-activation quantization, respectively, as
$\bm{W}^{\top}\bm{X}=$ $\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)},$ (4)
$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=$ $\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$
Finally, define the (normalized) blockwise output loss as
@@
We will approximate the layerwise loss as $\ell$ $\approx$ $\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise losses $\ell_{\left(i\right)}$ independently.
Optimal block-diagonal transforms. In the following, we will prove, under reasonable assumptions, that there exist closed-form optimal transforms for each block. These transforms are provably optimal for FP and asymptotically optimal for INT quantization. Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy
$\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=$ $\ d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top},$ (6)
$\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=$ $\ d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top},$
@@
respectively. Without loss of generality (Appendix Section˜A.1), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$. In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moments before the Cholesky decomposition. Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of
$\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}.$ (7)
Let $\bm{H}\in\bigl\{\pm d^{-\frac{1}{2}}\bigr\}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix. Then, we will show that the optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ can be constructed as
${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}=$ $\ \bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top},$ (8)
${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}=$ $\ \bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$
respectively. Note that [^1] The -$\top$ superscript means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$.
${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}={\bm{T}_{\mathrm{wush}}^{-\top}}_{\left(i\right)},$ (9)
@@
which can be easily verified by calculating ${\bm{T}_{\mathrm{xvsh}}}_{\left(i\right)}{\bm{T}_{\mathrm{wush}}^{\top}}_{\left(i\right)}=\mathbf{I}$, where $\mathbf{I}$ is the identity matrix.
@@
Remark. The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Equation˜8. Incidentally, this explains why it has been empirically observed to be an effective data-agnostic orthogonal transform.
1
21:Input:
32:weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, Hadamard $\bm{H}\in\bigl\{\pm d^{-\frac{1}{2}}\bigr\}^{d\times d}$, damping ratio $\lambda\in\mathbb{R}_{\geq 0}$, flag flagflag of whether to quantize with GPTQ
43:Output:
54:WUSH transform ${\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\tilde{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
65:
[…]
109:for $i\leftarrow d_{\mathrm{in}}/d$ to 11 do
[…]
1615:if flagflag then
1716:  $\tilde{\bm{W}}_{\left(i\right)}\leftarrow\operatorname{GPTQ}\left(\mathrm{weight}=\bar{\bm{W}},\mathrm{Hessian}=\bm{H}\bm{S}\bm{H}^{\top}\right)$
[…]
1917:  $\bm{Y}\leftarrow\bm{Y}-\tilde{\bm{W}}_{\left(i\right)}^{\top}{\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\bm{X}^{\prime}[j_{1}:j_{2},1:d_{\mathrm{in}}]$ $\triangleright$ inter-block GPTQ error propagation
2018:else
2119:  $\tilde{\bm{W}}_{\left(i\right)}\leftarrow q\left(\bar{\bm{W}}\right)$ $\triangleright$ RTN (round-to-nearest) fallback
2220:  $\bm{Y}\leftarrow\bm{Y}-\bar{\bm{W}}^{\top}{\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\bm{X}^{\prime}[j_{1}:j_{2},1:d_{\mathrm{in}}]$ $\triangleright$ blockwise triangular solve
2321:end if
2422:end for
Algorithm 1 Compute WUSH and Pre-Quantize Weights
GPTQ integration. In practice, the second-order information $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$ in Equation˜6 can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ. Therefore, it is a natural idea to integrate WUSH into GPTQ. However, GPTQ updates the weights iteratively via error propagation, while the WUSH transform is constructed from the second-order statistics of the updated weights. This coupling between weight updates and transform construction requires an interleaved computational schedule. Algorithm˜1 summarizes the procedure used to compute the blockwise [^2] The $\cdots\left[a:b,c:d\right]$ notation in algorithm selects block with row indices from $a$ to $b$ and column indices from $c$ to $d$, all inclusive. WUSH transforms and pre-quantized weights for a linear layer, using either GPTQ or RTN. The GPTQ error propagations are decomposed into intra-block and inter-block updates similar to its original paper (Frantar et al., 2023). For the intra-block updates, we apply the standard GPTQ subroutine to the current transformed weight block using the corresponding transformed Hessian. For inter-block propagation, we follow GPTQ’s usual blockwise updates, but we implement them using the equivalent formulation from Chen et al. (2025), which simplifies the procedure by avoiding explicit Hessian inversion. The model-level calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one. During inference, the forward pass of a linear layer is calculated as
$\sum_{i=1}^{d_{\mathrm{in}}/d}\tilde{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{wush}}}_{\left(i\right)}\tilde{\bm{X}}_{\left(i\right)}\right)$ (10)
with $\tilde{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ being the pre-quantized and transformed weight blocks and $\tilde{\bm{X}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{batch}}}$ being the new activation blocks instead of the calibration ones.